What is a in the equation 4(a +5) = 1-(a-(1-a))? The answer says 15/8 and I keep getting -15/4
Answer by rwm(914)

As it is written a=-3
4(a+5)=1-(a-(1-a))
remove last parentheses
4(a+5)=1-(a-1+a)
combine
4(a+5)=1-(2a-1)
remove last parentheses
4*(a+5)=1-2a+1
4*(a+5)=2-2a
combine
4(a+5)=2(1-a)
divide by 2
2(a+5)=1-a
remove last parentheses
2a+10=1-a
combine
3a=-9
divide by 3
a=-3
4(-3+5)=1-(-3-(1-(-3))
4*(2)=1-(-3-(1+3))
8=1-(-3-4)
8=1-(-7)
8=1+7
8=8
check